OpenStudy (trojanpoem):

Integrate from 0 to 1 tan^-1(x)/(1+x) dx

OpenStudy (trojanpoem):

\[\int\limits_{0}^{1}\frac{ \tan^{-1} x }{ 1+x } dx\]

satellite73 (satellite73):

you sure it is not \[\int\limits_{0}^{1}\frac{ \tan^{-1} x }{ 1+x^2 } dx\]

satellite73 (satellite73):

cause i seriously doubt you are going to find an anti derivative for \[\frac{\tan^{-1}(x)}{1+x}\]

OpenStudy (trojanpoem):

Nope, not 1+x^2 if it were 1+x^2 the final answer will be (tan^(-1)x)^2 / 2 ] from 0 to 1

satellite73 (satellite73):

where did the question come from? you are not going to find a closed form for your anti derivative

OpenStudy (trojanpoem):

.boredofstudies

OpenStudy (trojanpoem):

final answer pi/4 log2 - pi/8 log 2

OpenStudy (trojanpoem):

which is pi log2/8

satellite73 (satellite73):

wow how did you do it?

OpenStudy (trojanpoem):

:D wolfram xDDD

satellite73 (satellite73):

gotta love wolfram !

OpenStudy (trojanpoem):

No steps :/

OpenStudy (anonymous):

Here's a possible starting point. Call the integral \(I\). Take \(\arctan x=y\) (so that for \(|x|<\dfrac{\pi}{2}\) we have \(x=\tan y\)). This means \(\mathrm{d}x=\sec^2y\,\mathrm{d}y\). Now, \[I=\int_0^{\pi/4} \frac{y\sec^2y}{1+\tan y}\,\mathrm{d}y\] Integrating by parts, we could use \[\begin{matrix}u=y&&&\mathrm{d}v=\dfrac{\sec^2y}{1+\tan y}\,\mathrm{d}y\\[1ex] \mathrm{d}u=\mathrm{d}y&&&v=\ln(1+\tan y)\end{matrix}\]so that \[\begin{align*}I&=\bigg[y\ln(1+\tan y)\bigg]_{y=0}^{y=\pi/4}-\int_0^{\pi/4}\ln(1+\tan y)\,\mathrm{d}y\\[1ex] &=\frac{\pi\ln2}{4}-\int_0^{\pi/4}\ln(1+\tan y)\,\mathrm{d}y \end{align*}\]which gives you what you know is at least part of the answer. An interesting thing to note is that since the final result for the definite integral is expected to be \(\dfrac{\pi\ln2}{8}\), you could try showing that \(I\) is exactly the same as the remaining integral. That is, \[2I=\frac{\pi\ln2}{4}\implies I=\frac{\pi\ln2}{8}\] How to do this isn't immediately clear to me. Another possible route is to consider the series expansion of \(\ln(1+\xi)=\xi-\dfrac{\xi^2}{2}+\dfrac{\xi^3}{3}-\dfrac{\xi^4}{4}+\cdots\) with \(\xi=\tan y\), but I haven't done enough investigating to come up with anything conclusive.

OpenStudy (trojanpoem):

@anonymous For the remaining integral : use this property \[\int\limits_{0}^{\pi/4} \ln(1+tany) dy = \int\limits_{0}^{\pi/4} \ln(1+ \tan(\frac{ \pi }{ 4 } - y) dy\] \[\int\limits_{0}^{\pi/4} \ln(\frac{ 1 + tany }{ 1+ tany }+ \frac{ 1 - tany }{ 1+ tany}) dy\] \[\int\limits_{0}^{\pi/4} \ln(2) - \ln(1 + tany) dy\] \[2 I_{1} = \int\limits_{0}^{\pi/4} \ln(2) dy = [\ln(2) y ]_{0}^{\pi/4} = \ln(2)\frac{ \pi }{ 4 }\] \[I_{1} = \frac{ \ln2 \pi }{ 8 }\] Which adds up with your result to ln(2)pi/8 Nice work mate , keep it up.